OpenStudy (jennyrlz):

Express the complex number in trigonometric form. -4

OpenStudy (anonymous):

Do you have any answer choses? Or you have to write?

OpenStudy (jennyrlz):

i do one second. :)

OpenStudy (jennyrlz):

4(cos 270(degrees) + i sin 270 (degrees)) 4(cos 90(degrees)+ i sin 90(degrees)) 4(cos 0(degrees) + i sin 0(degrees)) 4(cos 180(degrees) + i sin 180(degrees))

OpenStudy (jennyrlz):

You still there?

OpenStudy (jennyrlz):

i think it is the last option but i am not sure >.<

OpenStudy (jennyrlz):

i would appreciate if you show your work :D

OpenStudy (anonymous):

\[-4=4\left( -1+0 i \right)=r(\cos \theta+i \sin \theta),\cos \theta=-1=\cos 180,\sin \theta=0=0,180\] \[\theta=180\] r=4 ?

OpenStudy (jennyrlz):

that's how i got 180 :D so was i right :)

OpenStudy (jennyrlz):

Thanks anyways! :)